OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). Choose one answer. a. y = 2x + 6 b. y = −one halfx + 3 c. y = −one halfx + 6 d. y = 2x + 3

OpenStudy (sachintha):

So can you work out?

OpenStudy (anonymous):

yeah

OpenStudy (rhystic):

So this is similar to the parallel problem except the rule for perpendicular lines is that the gradient/slopes are inverse of each other so if you had a slope of -3 the perpendicular line would have a slope of 3.

OpenStudy (sachintha):

Remember that this is not parallel. It's perpendicular to the line y = 2x+2.

OpenStudy (anonymous):

So the slope for the other line is -2?

OpenStudy (sachintha):

No

OpenStudy (anonymous):

or -1/2

OpenStudy (sachintha):

Yeah correct!

OpenStudy (anonymous):

So it is b or c?

OpenStudy (anonymous):

now i plug in the x?

OpenStudy (sachintha):

As previously check whether the y value is 3 by plugging x = 6.

OpenStudy (anonymous):

c?

OpenStudy (sachintha):

Can you show the calculation steps?

OpenStudy (anonymous):

y= -1/2+6

OpenStudy (anonymous):

-1/2x

OpenStudy (anonymous):

so 6 x -1/1 = 3

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

oh ok so y= -1/2x + 6 -1/2 x 6 = -3 -3 + 6 = 3 y=3

OpenStudy (rhystic):

here did you get the 6 from?

OpenStudy (anonymous):

the x ?

OpenStudy (sachintha):

I have to recheck too. :D

OpenStudy (rhystic):

youre math is a bit off. it should be b double check yotur work

OpenStudy (anonymous):

Im not sure where i went wrong ;/

OpenStudy (sachintha):

@Rhystic How can it be b?

OpenStudy (sachintha):

Kyle is right unless I am wrong.

OpenStudy (anonymous):

should i open up another thread and do another?

OpenStudy (sachintha):

Your wish. :)

OpenStudy (sachintha):

You can check these by plotting them online.

OpenStudy (rhystic):

no youre right my math was off sorry about that.